OpenStudy (aravindg):

find dy/dx y=sqrt x raised to root x raised again to root x until infinity find dy/dx y=sqrt x raised to root x raised again to root x until infinity @Mathematics

OpenStudy (aravindg):

...........

OpenStudy (jamesj):

If y = sqrt x ^ sqrt x ^ ... then ln y = ln ( sqrt x ^ ... ) now ln(a^b) = b ln a. So pull out the exponent, which is also y. Hence ln y = y ln (sqrt x) = 1/2 . y . ln x Now differentiate both sides with respect to y y'/y = 1/2 ( y' ln x - y / x) Now solve for y'